OpenStudy (anonymous):

Please help! Simplify: x^2-9/x^2-6x+9

OpenStudy (anonymous):

\[\frac{x^2-9}{x^2-6x+9}\] First, factor the top and the bottom into a product of binomials. This is usually the first step in these types of problems. \[\frac{(x+3)(x-3)}{(x+3)(x+3)}\] The x+3 term in the numerator will cancel with one in the denominator. \[\frac{x-3}{x+3}\]

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

No problem :)